Question 2 A new event has been proposed for the Winter Olympics. A sprinting athlete will leap onto a 20kg bobsled. The person and bobsled will then slide down an ice-covered ramp and into a spring with a carefully calibrated spring constant of 20.000 N/m. The athlete who compresses the spring the farthest wins the gold medal. 20° Which types of mechanical energy, Kinetic Energy (KE), Gravitational Poten Spring Potential Energy (Usp), are present in the athlete/sled system at the following times? a) ial Energy (Up), and Let h- Om be at the bottom of the ramp. Enter your answers on your scantron (problem numbers 9-12). A. KE 9. As the athlete jumps onto her sled 10. When the athlete at the bottom of the ramp, but before reaching the spring B. U D. KE and U E. KE and U 11. When the spring is halfway compressed 12. When the spring is fully compressecd b) What is conserved in the systern at the fellowing times? Enter your answers on your scantron (problem numbers 13-16). A. Mechanical Energy B. Momentum C. BothD. Nothing 13. As the athlete collides with her sled 14. As the athlete slides down the ramp 15. As the spring is being compressed 16. As the athlete accelerates during her sprint
media%2Fbd1%2Fbd17f0f5-3814-40de-9f46-77
Answer #1

Question 2)

9) D) KE and Ugravity

10) A) KE

11) E) KE and Usp

12) C) Uspring

13) B) Momentum

14) C) Both

15) A) Mechanical energy

16) Mechanical energy

c) m1 = 50 kg
m2 = 20 kg
u1 = 10 m/s
u2 = 0

Apply conservation of momentum


m1*u1 + m2*u2 = (m1 + m2)*v

50*10 + 0 = (50 + 20)*v

==> v = 500/70

= 7.14 m/s

d) Apply conservation of energy

(m1 + m2)*g*h + (1/2)*(m1 + m2)*v^2 = (1/2)*k*x^2

(50 + 20)*9.8*h + (1/2)*(50 + 20)*7.14^2 = (1/2)*20000*1.03^2

==> h = 12.9 m